### v1

# 22footnotetext: To whom correspondence should be addressed

Abstract
Irreducible Cartesian tensors (ICTs) play a crucial role in the design of equivariant graph neural networks, as well as in theoretical chemistry and chemical physics. Meanwhile, the design space of available linear operations on tensors that preserve symmetry presents a significant challenge. The ICT decomposition and a basis of this equivariant space are difficult to obtain for high-order tensors. After decades of research, we recently achieve an explicit ICT decomposition for $n=5$ (bonvicini2024irreducible) with factorial time/space complexity. This work, for the first time, obtains decomposition matrices for ICTs up to rank $n=9$ with reduced and affordable complexity, by constructing what we call path matrices. The path matrices are obtained via performing chain-like contraction with Clebsch-Gordan matrices following the parentage scheme. We prove and leverage that the concatenation of path matrices is an orthonormal change-of-basis matrix between the Cartesian tensor product space and the spherical direct sum spaces. Furthermore, we identify a complete orthogonal basis for the equivariant space, rather than a spanning set (pearce2023brauer), through this path matrices technique. We further extend our result to the arbitrary tensor product and direct sum spaces, enabling free design between different spaces while keeping symmetry. The Python code is available in the appendix where the $n=6,\dots,9$ ICT decomposition matrices are obtained in <0.1s, 0.5s, 1s, 3s, 11s, and 4m32s, respectively.
Keywords: irreducible Cartesian tensor, Clebsch-Gordan coefficient, parentage scheme, equivariant graph neural networks, spherical harmonics
1 Introduction
The deployment of equivariant neural networks in physics, chemistry, and robotics often requires certain equivariance or invariance to effectively incorporate inductive bias. Equivariant Graph Neural Networks (EGNNs) are well-suited for this purpose, as they can maintain equivariance while processing features within the graph (zaverkin2024higher; simeon2024tensornet; batzner20223). One of the most promising approach to constructing equivariant layers involves leveraging irreducible representations. Depending on the basis of the space, two distinct types of irreducible representations are utilized in different EGNNs: spherical (spherical tensors) (musaelian2023learning; batatia2022mace) and Cartesian irreducible representations (Irreducible Cartesian Tensors, ICTs) (zaverkin2024higher; schutt2021equivariant). Spherical tensors serve as the starting point for EGNN based on irreducible representations. However, they require Clebsch-Gordan transforms with significant computational overhead to derive desired irreducible representations from existing ones during each inference, making them computationally intensive. In contrast, ICTs are more efficient, as high-order representations can be easily computed using tensor products (specifically, outer products in EGNNs), making them more computationally friendly for most deep learning frameworks. The ICTs are obtained through the ICT decomposition matrices $H$,
$T^{(n;j;q)}=H^{(n;j;q)}T^{(n)},$ (1)
where $T^{(n;j;q)}$ is an ICT, $T^{(n)}$ is a given tensor, $n$ represents the rank of the tensor, $j$ is called the weight of the irreducible representation, and $q$ denotes the label (since the multiplicity for given $n$ and $j$ is sometimes greater than $1$). The decomposition matrices $H$ should be constructed such that $T^{(n;j;q)}$ are irreducible. The challenge lies in the fact that $H$ of rank $n>5$ are unknown. The recent development of EGNNs research benefits from the irreducible decomposition matrices $H$ obtained from the physics and chemistry communities. The decomposition matrices for $n=4$ have been established by andrews1982irreducible. However, only very recently have we obtained the decomposition matrices for $n=5$ (bonvicini2024irreducible). Due to these limitations, state-of-the-art EGNNs cannot freely manipulate high-order ICTs or can only manipulate those whose weight equals their rank (zaverkin2024higher).
It has been shown that $l=8$ irreducible representations can greatly benefit the model, but under the spherical basis (liao2023equiformerv2). Therefore, the matrices for high-order orthogonal ICT decomposition are of great importance. Meanwhile, another crucial question is which linear operations applied to the Cartesian tensors are equivariant, as these operations constitute the design space of EGNNs, in other words, what is the vector space of linear mappings $f$ such that
$\rho(g)f(T)=f(\rho(g)T)$ (2)
where $g\in G$, and $\rho$ is the matrix representation. This vector space is called an equivariant space, also denoted as ${\rm End}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$. More generally, for different input and output spaces, this equivariant space can be denoted as ${\rm Hom}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes l},\left(\mathbbm{R}^{3}\right)^{\otimes k}\right)$. pearce2023brauer leverages brauer1937algebras’s theorem to find the spanning set of the equivariant space. However, as the rank increases, the number of elements in the spanning set becomes greater than the dimension of the space. Thus, it is also essential to identify a basis to reduce the number of parameters required.
In this work, one of our contributions is to efficiently obtain ICT decomposition matrices for $6\leq n\leq 9$. Several previous methods (andrews1982irreducible; dincckal2013orthonormal; bonvicini2024irreducible) rely on evaluating the full permutation of $n$ (e.g., fundamental isotropic Cartesian tensors) and $2n$ (e.g., natural projection tensors) indices, which requires factorial complexity, making it impractical for application to high-rank tensors. bonvicini2024irreducible proposes a novel method to automate the procedure but does not reduce the complexity. The Reduced Row Echelon Form (RREF) algorithm is also required to run on an $n!\times n!$ matrix. Instead, we leverage the fact that the Cartesian tensor product space is equivalent to a direct product of spherical spaces, up to a change of basis. In spherical space, it is straightforward to identify the desired irreducible representations, as they are diagonalized; then, the inverted change of basis can bring us back to Cartesian space. The challenge lies in finding such a change-of-basis matrix. We prove that one can construct this matrix by concatenating the results of sequential contractions and the outer products of Clebsch-Gordan matrices (CG matrices) according to the parentage scheme (coope1965irreducible), which we refer to as path matrices. The change-of-basis matrix converts Cartesian tensors to the direct sum of spherical tensors.
If we treat tensors as vectors, we find that different spherical vectors are separated as distinct elements of the full-length vector. Thus, we can multiply it by a diagonalized matrix with only zeros and ones on the diagonals. This operation eliminates the other spherical vectors and retains only the desired one. The inverted change of basis can then convert the spherical direct product space back to the Cartesian tensor product space. The product of these three matrices forms the decomposition matrix, which decomposes the Cartesian tensor into the desired ICTs. This approach takes less than one second on CPUs to compute cases for $n=6$, which is already faster than the current decomposition for $n=5$. We also show that this process can be further accelerated. Specifically, we prove that path matrices have orthonormal columns up to multiplicative factors, which we leverage to demonstrate that a specific decomposition matrix can be generated via the product of the path matrix and its transpose.
Method Ranks Orthogonality Time/Space complexity
coope1965irreducible 3 No calculated by hand
andrews1982irreducible 4 No calculated by hand
dincckal2013orthonormal 3 Yes calculated by hand
bonvicini2024irreducible 5 No factorial
Ours 6-9 Yes exponential
Table 1: Comparison with previous works.
Another challenge this work addresses is the construction of bases for the equivariant spaces. With the support of Schur’s lemma, we prove that the bases are formed by matrices that: 1) decompose a vector into its ICT, and 2) map an ICT to another ICT in an isomorphic space (i.e., with the same weight). These matrices are easy to construct in the spherical direct sum space, as their elements lie on the diagonal. Then, our path matrices can change the basis from the spherical space to the Cartesian tensor product space. We further extend our conclusions to linear equivariant operations between different direct sum and tensor product spaces. Additionally, we demonstrate the capability to construct bases for such general equivariant spaces, including spherical tensor product spaces with different parities. This enables the efficient construction of equivariant layers between different spaces, thereby expanding the design space of EGNNs.
This work has three main contributions:
  • We propose a method to efficiently calculate high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$.
  • We characterize the bases of equivariant spaces and propose a method to efficiently obtain these bases, overcoming the redundancy in pearce2023brauer’s method.
  • We extend our approach to efficiently obtain bases for equivariant spaces operating between different spaces.
These contributions enable EGNNs to process high-rank ICTs and flexibly design equivariant layers between desired spaces. Furthermore, this work benefits the physics and chemistry communities by facilitating theoretical analysis based on high-rank ICTs.
The following contents are organized as follows. In Section 2, we recall some basic ideas of representation theory, spherical harmonics, tensors, tensor product spaces, and we also introduce the Clebsch-Gordan coefficients and the parentage scheme. In Section LABEL:sec:problem, we formulate our problems: obtaining high-order ICT decomposition matrices and bases for arbitrary tensor product spaces. In Section LABEL:sec:highorderict, we introduce our theory and method to obtain high-order ICT decomposition matrices. In Section LABEL:sec:basis, we present the theory and method to obtain bases of equivariant spaces where the input and output spaces are the same. In Section LABEL:sec:extension, we extend the theory to the equivariant design spaces where the input and output spaces are different, and show the method to obtain bases for such general spaces. In Section LABEL:sec:relatedworks, we review related works on designing EGNNs and calculating ICT decomposition matrices. In Section LABEL:sec:discussion, we discuss the relation between pearce2023brauer and this work, as well as the theory for $SO(n)$, $O(n)$, and $SU(n)$. We provide concluding remarks in Section LABEL:sec:conclusion, and discuss the potential social impact in Section LABEL:sec:boardersocialimpact. We introduce the reproducibility details in Section LABEL:sec:repro. The acknowledgments are presented after the main text. The code and visualizations are shown in the appendix.
2 Preliminaries
2.1 Representations and group equivariance
We first recall some basic facts. A thorough introduction to group theory can be found in inui2012group and brocker2013representations. We first focus on the $O(3)$ group (3D rotation and reflection), but we also extend our conclusions to $O(n)$, $SO(n)$, and $SU(n)$ in Section LABEL:sec:discussion. Consider a vector $v\in V$. The rotation and reflection operations should be able to act on it. Representation theory shows that $O(3)$ can be represented by some real matrices in the Cartesian space via a group homomorphism $\rho:G\to\mathbbm{R}^{3\times 3}$, which is called the natural representation. It holds that
$\rho(g_{1}*g_{2})=\rho(g_{1})\cdot\rho(g_{2})$
for any $g_{1}$, $g_{2}\in G$. The rotation $g$ acting on the vector $v$ can therefore be represented as a matrix multiplication $\rho(g)\cdot v$. Sometimes, one may also refer to the space of $v$ as a representation space. If the results of the group action remain in the same space, i.e., $\rho(g)v\in V$. In this sense, this place is also called a $G$-invariant space or subspace. To represent $O(3)$, the matrix $\rho(g)$ is constrained to be orthogonal. It is important to note that the co-domain of $\rho$ can differ, meaning it can also be represented by matrices of other dimensions. In this paper, we denote the co-domain of $\rho$ using a subscript when it should be made explicit, e.g., $\rho_{\mathbbm{R}^{3}}$. Now, it suffices to define equivariance. We claim that a function $f:X\to Y$ is equivariant if and only if
$\rho_{Y}f(x)=f(\rho_{X}(x)).$
We only consider $f$ as a linear function, and thus it is represented as a matrix. The set of all such $f$ constitutes a vector space. We sometimes denote this space as ${\rm End}_{O(3)}(X)$ if $X=Y$, and as ${\rm Hom}_{O(3)}(X,Y)$ otherwise. Meanwhile, we say that the function $f$ is invariant if $\rho_{Y}=1$. Since $X$ is not necessarily the same as $Y$, this actually makes finding the function $f$ relatively straightforward. Specifically, to fulfill
$\rho_{Y}(g)M=M\rho_{X}(g),$
we can simply let
$\rho_{Y}(g)=M\rho_{X}(g)M^{-1}.$
However, this approach is very time-consuming in the design of EGNNs. The matrix $M$ here is learnable in the networks. During training, we must repeatedly compute the inverse of $M$, which is computationally expensive. Even if we could afford this, the results of the $M_{i}v_{i}$ is not addable in general, as the co-domains of different $M_{i}$ vary. Moreover, we cannot guarantee that $M$ is actually invertible without adding restriction on the determinants.
Considering that linear combinations involve a tremendous number of inversion operations, this process becomes intractable. Therefore, we first focus on the simpler case where $X=Y$. We will later generalize this to the case where $X\neq Y$, but $Y$ is known and remains consistent across different $M$. When it is necessary to emphasize the change in space, we refer to a linear mapping $M$ as $G$-$(X,Y)$-equivariant.
2.2 Irreducible representations and spherical spaces
We then introduce the definition of reducible and irreducible representation, in a matrix view:
Definition 1
A group homomorphism $\rho\in{\rm Hom}_{G}(\mathbbm{R}^{k})$ is said to be reducible, if there is an invertible matrix $P$, such that for every $g\in G$, we have
$P\rho(g)P^{-1}={\rm diag}\big{(}\rho_{1}(g),\rho_{2}(g)\big{)}$
for some non-trivial $\rho_{1}$ and $\rho_{2}$.
The $3\times 3$ rotation and reflection matrices are irreducible. More generally,