### v2

# 22footnotetext: To whom correspondence should be addressed

Abstract
Irreducible Cartesian tensors (ICTs) play a crucial role in the design of equivariant graph neural networks, as well as in theoretical chemistry and chemical physics. Meanwhile, the design space of available linear operations on tensors that preserve symmetry presents a significant challenge. The ICT decomposition and a basis of this equivariant space are difficult to obtain for high-order tensors. After decades of research, we recently achieve an explicit ICT decomposition for $n=5$ (bonvicini2024irreducible) with factorial time/space complexity. This work, for the first time, obtains decomposition matrices for ICTs up to rank $n=9$ with reduced and affordable complexity, by constructing what we call path matrices. The path matrices are obtained via performing chain-like contraction with Clebsch-Gordan matrices following the parentage scheme. We prove and leverage that the concatenation of path matrices is an orthonormal change-of-basis matrix between the Cartesian tensor product space and the spherical direct sum spaces. Furthermore, we identify a complete orthogonal basis for the equivariant space, rather than a spanning set (pearce2023brauer), through this path matrices technique. We further extend our result to the arbitrary tensor product and direct sum spaces, enabling free design between different spaces while keeping symmetry. The Python code is available in https://github.com/ShihaoShao-GH/ICT-decomposition-and-equivariant-bases where the $n=6,\dots,9$ ICT decomposition matrices are obtained in 1s, 3s, 11s, and 4m32s, respectively.
Keywords: irreducible Cartesian tensor, Clebsch-Gordan coefficient, parentage scheme, equivariant graph neural networks, spherical harmonics
1 Introduction
The deployment of equivariant neural networks in physics, chemistry, and robotics often requires certain equivariance or invariance to effectively incorporate inductive bias. Equivariant Graph Neural Networks (EGNNs) are well-suited for this purpose, as they can maintain equivariance while processing features within the graph (zaverkin2024higher; simeon2024tensornet; batzner20223). One of the most promising approach to constructing equivariant layers involves leveraging irreducible representations. Depending on the basis of the space, two distinct types of irreducible representations are utilized in different EGNNs: spherical (spherical tensors) (musaelian2023learning; batatia2022mace) and Cartesian irreducible representations (Irreducible Cartesian Tensors, ICTs) (zaverkin2024higher; schutt2021equivariant) for $O(3)$. Spherical tensors serve as the starting point for EGNN based on irreducible representations. However, they require Clebsch-Gordan transforms with significant computational overhead to derive desired irreducible representations from existing ones during each inference, making them computationally intensive. In contrast, ICTs are more efficient, as high-order representations can be easily computed using tensor products (specifically, outer products in EGNNs), making them more computationally friendly for most deep learning frameworks. The ICTs are obtained through the ICT decomposition matrices $H$,
${\rm vec}\left(T^{(n;j;q)}\right)=H^{(n;j;q)}{\rm vec}\left(T^{(n)}\right),$ (1)
where $T^{(n;j;q)}$ is an ICT, ${\rm vec}$ is the vectorization operation via flattening the tensor, $T^{(n)}$ is a given tensor, $n$ represents the rank of the tensor, and $j$ is called the weight (the order of irreducible representations, which is distinct from the learnable weights in neural networks) of the irreducible representation, and $q$ denotes the label (since the multiplicity for given $n$ and $j$ is sometimes greater than $1$). The decomposition matrices $H^{(n;j;q)}$ should be constructed such that $T^{(n;j;q)}$ are irreducible. A formal definition is introduced in Section LABEL:sec:ictdecomp. The challenge lies in the fact that $H$ of rank $n>5$ are unknown. The recent development of EGNNs research benefits from the irreducible decomposition matrices $H$ obtained from the physics and chemistry communities. The decomposition matrices for $n=4$ have been established by andrews1982irreducible. However, only very recently have we obtained the decomposition matrices for $n=5$ (bonvicini2024irreducible). Due to these limitations, state-of-the-art EGNNs cannot freely manipulate high-order ICTs or can only manipulate those whose weight equals their rank (zaverkin2024higher).
It has been shown that $l=8$ irreducible representations can greatly benefit the model, but under the spherical basis (liao2023equiformerv2). Therefore, the matrices for high-order orthogonal ICT decomposition are of great importance. Meanwhile, another crucial question is which linear operations applied to the Cartesian tensors are equivariant, as these operations constitute the design space of EGNNs, in other words, what is the vector space of linear mappings $f$ such that
$\rho(g)f\left({\rm vec}(T)\right)=f\left(\rho(g){\rm vec}(T)\right)$ (2)
where $g\in G$, and $\rho$ is the matrix representation. This vector space is called an equivariant space, also denoted as ${\rm End}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$. More generally, for different input and output spaces, this equivariant space can be denoted as ${\rm Hom}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes l},\left(\mathbbm{R}^{3}\right)^{\otimes k}\right)$. pearce2023brauer leverages brauer1937algebras’s theorem to find the spanning set of the equivariant space. However, as the rank increases, the number of elements in the spanning set becomes greater than the dimension of the space. Thus, it is also essential to identify a basis to reduce the number of parameters required.
In this work, one of our contributions is to efficiently obtain ICT decomposition matrices for $6\leq n\leq 9$. Several previous methods (andrews1982irreducible; dincckal2013orthonormal; bonvicini2024irreducible) rely on evaluating the full permutation of $n$ (e.g., fundamental isotropic Cartesian tensors) and $2n$ (e.g., natural projection tensors) indices, which requires factorial complexity, making it impractical for application to high-rank tensors. bonvicini2024irreducible proposes a novel method to automate the procedure but does not reduce the complexity. The Reduced Row Echelon Form (RREF) algorithm is also required to run on an $n!\times n!$ matrix. Instead, we leverage the fact that the Cartesian tensor product space is equivalent to a direct product of spherical spaces, up to a change of basis. In spherical space, it is straightforward to identify the desired irreducible representations, as they are diagonalized; then, the inverted change of basis can bring us back to Cartesian space. The challenge lies in finding such a change-of-basis matrix. We prove that one can construct this matrix by concatenating the results of sequential contractions and the outer products of Clebsch-Gordan matrices (CG matrices) according to the parentage scheme (coope1965irreducible), which we refer to as path matrices. The change-of-basis matrix converts Cartesian tensors to the direct sum of spherical tensors.
If we treat tensors as vectors, we find that different spherical vectors are separated as distinct elements of the full-length vector. Thus, we can multiply it by a diagonalized matrix with only zeros and ones on the diagonals. This operation eliminates the other spherical vectors and retains only the desired one. The inverted change of basis can then convert the spherical direct product space back to the Cartesian tensor product space. The product of these three matrices forms the decomposition matrix, which decomposes the Cartesian tensor into the desired ICTs. This approach takes less than one second on CPUs to compute cases for $n=6$, which is already faster than the current decomposition for $n=5$. We also show that this process can be further accelerated. Specifically, we prove that path matrices have orthonormal columns up to multiplicative factors, which we leverage to demonstrate that a specific decomposition matrix can be generated via the product of the path matrix and its transpose.
Method Rank Orthogonality Time/Space complexity
coope1965irreducible 3 No calculated by hand
andrews1982irreducible 4 No calculated by hand
dincckal2013orthonormal 3 Yes calculated by hand
bonvicini2024irreducible 5 No factorial
Ours 6-9 Yes exponential
Table 1: Comparison with previous works. The complexity is discussed in terms of rank. The complexity of this work will be polynomial if we focus on the dimensions.
Another challenge this work addresses is the construction of bases for the equivariant spaces. With the support of Schur’s lemma, we prove that the bases are formed by matrices that: 1) decompose a vector into its ICT, and 2) map an ICT to another ICT in an isomorphic space (i.e., with the same weight). These matrices are easy to construct in the spherical direct sum space, as their elements lie on the diagonal. Then, our path matrices can change the basis from the spherical space to the Cartesian tensor product space. We further extend our conclusions to linear equivariant operations between different direct sum and tensor product spaces. Additionally, we demonstrate the capability to construct bases for such general equivariant spaces, including spherical tensor product spaces with different parities. This enables the efficient construction of equivariant layers between different spaces, thereby expanding the design space of EGNNs.
This work has four main contributions:
  • We propose a method to efficiently calculate high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$, compared to predecessors, as shown in Table 1.
  • We propose a general parentage scheme to extend ICT decomposition into a generalized ICT decomposition applicable to arbitrary tensor product spaces.
  • We characterize the bases of equivariant spaces and propose a method to efficiently obtain these bases, overcoming the redundancy in pearce2023brauer’s method.
  • We extend our approach to efficiently obtain bases for equivariant spaces operating between different spaces.
These contributions enable EGNNs to process high-rank ICTs and flexibly design equivariant layers between desired spaces. Furthermore, this work benefits the physics and chemistry communities by facilitating theoretical analysis based on high-rank ICTs.
The following contents are organized as follows. In Section 1, we recall some basic ideas of representation theory, spherical harmonics, tensors, tensor product spaces, and we also introduce the Clebsch-Gordan coefficients and the parentage scheme. In Section LABEL:sec:problem, we formulate our problems: obtaining high-order ICT decomposition matrices and bases for arbitrary tensor product spaces. In Section LABEL:sec:highorderict, we introduce our theory and method to obtain high-order ICT decomposition matrices. In Section LABEL:sec:basis, we present the theory and method to obtain bases of equivariant spaces where the input and output spaces are the same. In Section LABEL:sec:extension, we extend the theory to the equivariant design spaces where the input and output spaces are different, and show the method to obtain bases for such general spaces. In Section LABEL:sec:relatedworks, we review related works on designing EGNNs and calculating ICT decomposition matrices. In Section LABEL:sec:discussion, we discuss the relation between pearce2023brauer and this work, as well as the theory for $SO(n)$, $O(n)$, and $SU(n)$. We provide concluding remarks in Section LABEL:sec:conclusion, and discuss the potential social impact in Section LABEL:sec:boardersocialimpact. We introduce the reproducibility details in Section LABEL:sec:repro. The acknowledgments are presented after the main text. The visualizations are shown in the appendix. A list of mathematical symbols is provided in Table LABEL:tab:symbols, and a mind map of this work is shown in Figure 1.
Table 2: Mathematical symbols and their meanings
Symbol Meaning Symbol Meaning
$\lVert\cdot\rVert_{col}$ The $L_{2}$-norm of an arbitrary column of $\cdot$ $G$ An arbitrary group
$SO(n)$ Special orthogonal group on $n$-dimensional space $SU(n)$ Special unitary group on $n$-dimensional space
$O(n)$ Orthogonal group on $n$-dimensional space $U(n)$ Unitary group on $n$-dimensional space
$\mathcal{S}$ A vector space $\rho_{\mathcal{S}}$ A group representation
$\mathcal{S}^{\otimes n}$ A rank $n$ tensor product space from $\mathcal{S}$ $N^{(n;j)}$ The multiplicity of the ICT of rank $n$ and weight $j$
$T$ An arbitrary tensor $G\to\mathcal{S}$ $T^{(n)}$ A Cartesian tensor of rank $n$
$\hat{T}^{(n;j;q;p)}$ An ICT of rank $n$, weight $j$, index $q$, and parity $p$ (sometimes omitted) $Q^{(l;p)}_{\mathcal{S}}$ The set of all possible paths to decompose a tensor in space $\mathcal{S}$
$H^{(n;j;q)}$ An ICT decomposition matrix to obtain ICTs of rank $n$, weight $j$, and index $q$ $\hat{H}^{(\mathcal{S};j;q)}$ A general ICT decomposition matrix of space $\mathcal{S}$, weight $j$ and index $q$
$C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}$ A CG coefficient with input weights $l_{1}$, $l_{2}$, output weight $l_{o}$, and the corresponding indices $m_{1}$, $m_{2}$, and $m_{o}$ (the angular-momentum quantum numbers in physics)
$C^{l_{1}l_{2}l_{o}}$ A CG tensor with input weights $l_{1}$, $l_{2}$, and output weight $l_{o}$ $\hat{C}^{l_{1}l_{2}l_{o}}$ A CG matrix by merging the first two dimensions of the original CG tensor
$P^{(path)}$ A path matrix generated by $path$, e.g., $(1\to 3\to 2)$ $\hat{P}^{(path)}$ A normalized path matrix generated by $path$
$D^{(\mathcal{S}\to\mathcal{S}^{\prime})}$ A change-of-basis matrix from space $\mathcal{S}$ to $\mathcal{S}^{\prime}$ $v^{(l;p)}$ A representation weight $l$ and parity $p$ irreducible representation acts on
${\rm Hom}_{G}$ The set of $G$-equivariant maps from one space to another ${\rm End}_{G}$ The set of $G$-equivariant maps from one space to itself
irr Returns all weights of irreducible representations in a decomposition ${\rm dim}$ The number of dimensions of a space
$(i_{1}\overset{k_{1}}{\to}i_{2}$ $\overset{k_{2}}{\to}\dots)$ A path where $k_{*}$ are bridge numbers (omitted if all being $1$) $(l,p)$ An irreducible representation space of weight $l$ and parity $p$
$\otimes$ Tensor product $\oplus$ Direct sum
$\odot$ Contraction ##\## Number of elements in a set
2 Preliminaries
2.1 Clebsch-Gordan coefficients
A review on irreducible representations and spherical spaces is in the appendix. The term weights in this paper refer to the order of irreducible representations. An important tool utilized in our work is the Clebsch-Gordan coefficients (CG coefficients). They are real numbers, and describe how to couple vectors from two spherical spaces into a new one. There are some rules for this coupling. First, given two vectors in weights $l_{1}$ and $l_{2}$ spherical spaces, the weights of the resulting vectors can only lie in $|l_{1}-l_{2}|\leq l_{o}\leq l_{1}+l_{2}$. For each index $m_{o}$ of the obtained vector $v_{l_{o}}\in(l=l_{o})$, it is calculated by homogeneous quadratic polynomials. Specifically, we have
$v^{(l_{o})}_{m_{o}}=\sum_{m_{o}=m_{1}+m_{2}}C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}v_{m_{1}}^{(l_{1})}v_{m_{2}}^{(l_{2})},$ (3)
and if we consider O(n) instead of SO(n), parity $p$ should also be considered. The parity is a property additional to the rank and weight, being $1$ or $-1$. Having $-1$ parity means that the representation will be inverse when the space is inverse, and vice versa. For example, the natural representations of O(n) always have $p=-1$. We will introduce the case that $p=1$ in the next section. Taking the parity into consideration, we can guarantee O(n) equivariance, and equation (3) becomes
$v^{(l_{o};p_{o})}_{m_{o}}=\mathbbm{1}_{(p_{o}=p_{1}p_{2})}\sum_{m_{o}=m_{1}+m_{2}}C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}v_{m_{1}}^{(l_{1};p_{1})}v_{m_{2}}^{(l_{2};p_{2})},$
where $\mathbbm{1}_{(exp)}$ is the indicator function, outputting $1$ if $exp$ holds true, or 00 otherwise. Our discussion will be the same no matter we consider $O(n)$ or $SO(n)$. But since Section LABEL:sec:extension, we need to take care of the parity which will play some roles in the further discussion. The CG coefficients are highly sparse, since $C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}=0$ if $m_{o}\neq m_{1}+m_{2}$. The coefficients are identical to Wigner-3j symbols up to multiplicative factors, so we will adopt Wigner-3j for the implementation. The evaluation of CG coefficients is nearly $\mathcal{O}(1)$ (if we do not consider the few calculations for the closed-form formula, which is cheap enough). To obtain the coefficients of $O(3)/SO(3)$, one often leverages the relation between $SU(2)$ and $SO(3)$: $SU(2)$ is a double cover of $SO(3)$, so the CG matrices of $SU(2)$ and $SO(3)$ are same up to a change-of-basis, from complex to real. The closed-form formulae for $SU(2)$ CG coefficients are already known. Thus, this process can be efficiently evaluated without obvious overhead.
Figure 1: The mind map of our method. The blocks in orange color are aligned with the three contributions of this work.
2.2 Tensors and tensor product spaces
To formally define a tensor, we need to consider vector space $V$ and covector space $V^{*}$. A tensor product is a bilinear map $\otimes:\mathcal{S}\times\mathcal{S}^{\prime}\to\mathcal{S}\otimes\mathcal{S}^{\prime}$. The basis of $\mathcal{S}\otimes\mathcal{S}^{\prime}$ is the set of $\hat{x}_{i}\otimes\hat{y}_{j}\in\mathcal{S}\otimes\mathcal{S}^{\prime}$, where $\hat{x}_{i}$ and $\hat{y}_{j}$ are basis vectors of $\mathcal{S}$ and $\mathcal{S}^{\prime}$, respectively. That is why one implements outer product as tensor product, as it meets the definition of tensor product. A tensor $T:V^{\otimes r}\otimes V^{*\otimes s}\to F$, where $F$ is some field, is a mathematical object that converts things between different tensor product spaces. For example, the representation $\rho(g)$ is a $(r=1,s=1)$-tensor that receives a vector $v$ which is a $(r=1,s=0)$-tensor, and output another $(r=1,s=0)$-tensor $\rho(g)v$, as $r$ and $s$ vanish correspondingly. Tensor product can help us build up the ranks of tensors, e.g., $V\otimes V\otimes V$ is a $(r=3,s=0)$ space. The representation space is accordingly $(r=3,s=3)$. Conversely, contraction can reduce the ranks of tensors. We express these operations via some index notations. For example, $T_{i_{1}i_{2}}=T^{\prime}_{i_{1}}T^{\prime\prime}_{i_{2}}$ is a tensor product, and $T_{i_{1}}=\sum_{j_{1}}T^{\prime}_{i_{1}j_{1}}T^{\prime\prime}_{j_{1}}$ is a contraction. Obviously, such operations can be mixed together, and we often omit the index of the output tensors and agree $i_{k}$ will be placed at the $k$th index of the output tensor by default. We also agree that $j_{1}$ and $j_{2}$ will vanish due to the contraction. We will not use Einstein summation convention in this work to keep the conventions of machine learning community, and to avoid ambiguous during derivations. One should note that contractions can be expressed in term of matrix representation. Consider $T=\sum_{j_{1},j_{2}}T^{\prime}_{i_{1}j_{1}j_{2}}T^{\prime\prime}_{j_{1}j_{2}}$, we have
$\begin{bmatrix}T_{1}\\ T_{2}\end{bmatrix}=\begin{bmatrix}T^{\prime}_{111}&T^{\prime}_{112}&T^{\prime}_{121}&T^{\prime}_{122}\\ T^{\prime}_{211}&T^{\prime}_{212}&T^{\prime}_{221}&T^{\prime}_{222}\end{bmatrix}\begin{bmatrix}T^{\prime\prime}_{11}\\ T^{\prime\prime}_{12}\\ T^{\prime\prime}_{21}\\ T^{\prime\prime}_{22}\end{bmatrix}.$
It is useful when we need to obtain the inverse of some linear operations. Also, contractions and tensor products are both $O(n)$-equivariant.
2.3 CG matrices and the parentage scheme
Figure 2: The parentage scheme. The numbers in the circle represent multiplicity. The sum of each column equals equation (LABEL:eq:multiplicity).
The $O(3)$ group has its representations in tensor product space $(\mathbbm{R}^{3})^{\otimes n}$, which are some $3^{n}\times 3^{n}$ matrices, or tensor product of $n$ $(r=1,s=1)$-tensors, which depends on one’s view. If we interpret them as matrices, then the objects they act on are simply $3^{n}$-dimensional vectors. The tensor product spaces $(\mathbbm{R}^{3})^{\otimes n}$ for $n>1$ are generally reducible. Recall CG coefficients, two irreducible spherical vectors $v$ and $v^{\prime}$ coupling can also be viewed as that they first form a rank 2 tensor, and then decompose into different new spherical vectors. These operations can also be viewed as a product of matrices and vectors. In this context, the CG coefficients $\hat{C}^{l_{1}l_{2}l_{o}}$ form a $(2l_{1}+1)(2l_{2}+1)\times(2l_{o}+1)$ matrix, which we call a CG matrix. If we reshape it to a $(2l_{1}+1)\times(2l_{2}+1)\times(2l_{o}+1)$ tensor, then we instead call it an $(l_{1},l_{2},l_{o})$-tensor $C^{l_{1}l_{2}l_{o}}$. Here are some important properties,
Lemma 1 (Orthogonality)
Given an $(l_{1},l_{2},l_{o})$-CG tensor $C^{l_{1}l_{2}l_{o}}$, we have
$\sum_{j_{1},j_{2}}C^{l_{1}l_{2}l_{o}}_{j_{1}j_{2}k}C^{l_{1}l_{2}l_{o}}_{j_{1}j_{2}k^{\prime}}=\sum_{j_{1},j_{2}}C^{l_{1}l_{2}l_{o}}_{j_{1}kj_{2}}C^{l_{1}l_{2}l_{o}}_{j_{1}k^{\prime}j_{2}}=\sum_{j_{1},j_{2}}C^{l_{1}l_{2}l_{o}}_{kj_{1}j_{2}}C^{l_{1}l_{2}l_{o}}_{k^{\prime}j_{1}j_{2}}=0$
for $k\neq k^{\prime}$. More generally, different CG tensors with the same $l_{1}$ and $l_{2}$, $l_{1}$ and $l_{3}$, or $l_{2}$ and $l_{3}$ are also orthogonal in this way.