### v6

# 22footnotetext: To whom correspondence should be addressed

Abstract
Irreducible Cartesian tensors (ICTs) play a crucial role in the design of equivariant graph neural networks, as well as in theoretical chemistry and chemical physics. Meanwhile, the design space of available linear operations on tensors that preserve symmetry presents a significant challenge. The ICT decomposition and a basis of this equivariant space are difficult to obtain for high-rank tensors. After decades of research, bonvicini2024irreducible recently achieves an explicit ICT decomposition for $n=5$ with factorial time/space complexity. In this work we, for the first time, obtains decomposition matrices for ICTs up to rank $n=9$ with reduced and affordable complexity, by constructing what we call path matrices. The path matrices are obtained via performing chain-like contractions with Clebsch-Gordan matrices following the parentage scheme. We prove and leverage that the concatenation of path matrices is an orthonormal change-of-basis matrix between the Cartesian tensor product space and the spherical direct sum spaces. Furthermore, we identify a complete orthogonal basis for the equivariant space, rather than a spanning set (pearce2023brauer), through this path matrices technique. To the best of our knowledge, this is also the first analytic, rather than numerical, method for theoretically obtaining arbitrary rank orthogonal ICT decomposition matrices and orthogonal equivariant bases. We further extend our result to the arbitrary tensor product and direct sum spaces, enabling free design between different spaces while keeping symmetry. The Python code is available at https://github.com/ShihaoShao-GH/ICT-decomposition-and-equivariant-bases, where the $n=6,\dots,9$ ICT decomposition matrices are obtained in 1s, 3s, 11s, and 4m32s on 28-cores Intel® Xeon® Gold 6330 CPU @ 2.00GHz, respectively.
Keywords: irreducible Cartesian tensor decomposition, Clebsch-Gordan coefficient, parentage scheme, equivariant graph neural network, equivariant linear mapping
1 Introduction
The deployment of equivariant neural networks in physics, chemistry, and robotics often requires certain equivariance or invariance to effectively incorporate inductive bias. Equivariant Graph Neural Networks (EGNNs) are well-suited for this purpose, as they can maintain equivariance while processing features within the graph (zaverkin2024higher; simeon2024tensornet; batzner20223). One of the most promising approach to constructing equivariant layers involves leveraging irreducible representations. Depending on the basis of the space, two distinct types of irreducible representations are utilized in different EGNNs: spherical (spherical tensors) (musaelian2023learning; batatia2022mace) and Cartesian irreducible representations (Irreducible Cartesian Tensors, ICTs) (zaverkin2024higher; schutt2021equivariant) for $O(3)$. Spherical tensors serve as the starting point for EGNNs based on irreducible representations. However, they require Clebsch-Gordan transforms with significant computational overhead to derive desired irreducible representations from existing ones during each inference, making them computationally intensive. In contrast, ICTs are more efficient, as high-rank representations can be easily computed using tensor products (specifically, outer products in EGNNs) and contractions, making them more computationally friendly for most deep learning frameworks. The ICTs are obtained through the ICT decomposition matrices $H$,
${\rm vec}\left(\hat{T}^{(n;l;q)}\right)=H^{(n;l;q)}{\rm vec}\left(T^{(n)}\right),$ (1)
where $\hat{T}^{(n;l;q)}$ is an ICT, ${\rm vec}$ is the vectorization operation via flattening the tensor, $T^{(n)}$ is a given tensor, $n$ represents the rank of the tensor, and $l$ is called the weight (the order of irreducible representations, which is distinct from the learnable weights in neural networks) of the irreducible representation, and $q$ denotes the label (since the multiplicity for given $n$ and $l$ is sometimes greater than $1$). The decomposition matrices $H^{(n;l;q)}$ should be constructed such that $\hat{T}^{(n;l;q)}$ are irreducible. A formal definition is introduced in Section LABEL:sec:ictdecomp. The challenge lies in the fact that $H^{(n;l;q)}$ of rank $n>5$ are unknown. The recent development of EGNNs research benefits from the irreducible decomposition matrices $H$ obtained from the physics and chemistry communities. The decomposition matrices for $n=4$ have been established by andrews1982irreducible. However, only very recently has bonvicini2024irreducible obtained the decomposition matrices for $n=5$. Due to these limitations, state-of-the-art EGNNs cannot freely manipulate high-rank ICTs or can only manipulate those whose weights equal their ranks (zaverkin2024higher).
It has been shown that $l=8$ irreducible representations can greatly benefit the model, but under the spherical basis (liao2023equiformerv2). Therefore, the matrices for high-rank orthogonal ICT decomposition are of great importance. Meanwhile, another crucial question is which linear operations applied to the Cartesian tensors are equivariant, as these operations constitute the design space of EGNNs, in other words, what is the vector space of linear mappings $f$ such that
$\rho(g)f\left({\rm vec}(T)\right)=f\left(\rho(g){\rm vec}(T)\right)$ (2)
where $g\in G$, $T\in\left(\mathbbm{R}^{3}\right)^{\otimes n}$, and $\rho:G\to GL\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$ is the matrix representation. This vector space is called an equivariant space, also denoted as ${\rm End}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$. More generally, for different input and output spaces, this equivariant space can be denoted as ${\rm Hom}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes l},\left(\mathbbm{R}^{3}\right)^{\otimes k}\right)$. pearce2023brauer leverages brauer1937algebras’s theorem to find the spanning set of the equivariant space. However, as the rank increases, the number of elements in the spanning set becomes greater than the dimension of the space. Thus, it is also essential to identify a basis to reduce the number of parameters required.
In this work, one of our contributions is the proposal of a non-numerical method for efficiently obtaining ICT decomposition matrices, through which we achieve $6\leq n\leq 9$ in an affordable setting. Several previous methods (andrews1982irreducible; dincckal2013orthonormal; bonvicini2024irreducible) rely on evaluating the full permutation of $n$ (e.g., fundamental isotropic Cartesian tensors) and $2n$ (e.g., natural projection tensors) indices, which requires factorial complexity, making it impractical for application to high-rank tensors. bonvicini2024irreducible proposes a novel method to automate the procedure but does not reduce the complexity. The Reduced Row Echelon Form (RREF) algorithm is also required to run on an $n!\times n!$ matrix. Instead, we leverage the fact that the Cartesian tensor product space is equivalent to a direct product of spherical spaces, up to a change of basis. In spherical space, it is straightforward to identify the desired irreducible representations, as they are diagonalized; then, the inverse change-of-basis can bring us back to Cartesian space. The challenge lies in finding such a change-of-basis matrix. We prove that one can construct this matrix by concatenating the results of sequential contractions and the outer products of Clebsch-Gordan matrices (CG matrices) according to the parentage scheme (coope1965irreducible), which we refer to as path matrices. The change-of-basis matrix converts Cartesian tensors to the direct sum of spherical tensors.
If we treat tensors as vectors, we find that different spherical vectors are separated as distinct elements of the full-length vector. Thus, we can multiply it by a diagonal matrix with only zeros and ones on the diagonal. This operation eliminates the other spherical vectors and retains only the desired ones. The inverted change of basis can then convert the spherical direct product space back to the Cartesian tensor product space. The product of these three matrices—the change-of-basis matrix, the diagonal matrix, and the inverse change-of-basis—forms the decomposition matrix, which decomposes the Cartesian tensor into the desired ICTs. This approach takes less than one second on 28-cores Intel® Xeon® Gold 6330 CPU @ 2.00GHz to compute cases for $n=6$, which is already much faster than the current decomposition for $n=5$ by bonvicini2024irreducible. We also show that this process can be further accelerated. Specifically, we prove that path matrices have orthonormal columns up to multiplicative factors, which we leverage to demonstrate that decomposition matrices can be generated via the product of the path matrices and their transpose, with no inverse operation involved.
Method Rank Orthogonality Time and Space Complexity
w.r.t. rank w.r.t. dimension
coope1965irreducible $\leq 3$ No calculated by hand
andrews1982irreducible $\leq 4$ No calculated by hand
dincckal2013orthonormal $\leq 3$ Yes calculated by hand
bonvicini2024irreducible $\leq 5$ No
factorial  / sub-exponential
Ours $\leq 9$ Yes
exponential  / polynomial
Table 1: Comparison with previous works. The time and space complexities are with respect to rank and dimension, respectively.
Another challenge this work addresses is the construction of bases for the equivariant spaces. With the support of Schur’s lemma, we prove that the bases are formed by matrices that: 1) decompose a vector into its ICT, and 2) map an ICT to another ICT in an isomorphic space (i.e., with the same weight). These matrices are easy to construct in the spherical direct sum space, as their elements lie on the diagonal. Then, our path matrices can change the basis from the spherical space to the Cartesian tensor product space. We further extend our conclusions to linear equivariant operations between different direct sum and tensor product spaces. We demonstrate the capability to construct bases for such general equivariant spaces, including spherical tensor product spaces with different parities. This enables the efficient construction of equivariant layers between different spaces, thereby expanding the design space of EGNNs.
In summary, this work has five main contributions:
  • We propose a non-numerical method to efficiently obtain high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$, significantly improving upon the current limit of $n=5$, as shown in Table 1.
  • We propose a general parentage scheme to extend ICT decomposition into a generalized ICT decomposition applicable to arbitrary tensor product spaces.
  • We propose to directly obtain orthogonal bases of equivariant spaces, whereas the method proposed by pearce2023brauer yields a spanning set, requiring an expensive post-processing step to extract a basis that is not orthogonal in general.
  • To the best of our knowledge, this is the first analytic, rather than numerical, method for theoretically obtaining arbitrary rank orthogonal ICT decomposition matrices and orthogonal equivariant bases.
  • We extend our approach to efficiently obtain bases of equivariant spaces mapping between different spaces, including spherical tensor product spaces.
These contributions enable EGNNs to process high-rank ICTs and flexibly design equivariant layers between desired spaces. Furthermore, this work benefits the physics and chemistry communities by facilitating theoretical analysis based on high-rank ICTs.
Table 2: Mathematical symbols and their meanings
Symbol Meaning Symbol Meaning
$\lVert\cdot\rVert_{col}$ The $L_{2}$-norm of an arbitrary column of $\cdot$ $G$ An arbitrary group
$SO(n)$ Special orthogonal group on $n$-dimensional space $SU(n)$ Special unitary group on $n$-dimensional space
$O(n)$ Orthogonal group on $n$-dimensional space $U(n)$ Unitary group on $n$-dimensional space
$\mathcal{V}$ A vector space $\rho_{\mathcal{V}}$ A group representation
$\mathcal{V}^{\otimes n}$ A rank $n$ tensor product space from $\mathcal{V}$ $N^{(n;j)}$ The multiplicity of the ICT of rank $n$ and weight $j$
$T$ An arbitrary tensor $G\to\mathcal{V}$ $T^{(n)}$ A Cartesian tensor of rank $n$
$\hat{T}^{(n;j;q;p)}$ An ICT of rank $n$, weight $j$, index $q$, and parity $p$ (sometimes omitted) $Q^{(l;p)}_{\mathcal{V}}$ The set of all possible paths to decompose a tensor in space $\mathcal{V}$
$H^{(n;j;q)}$ An ICT decomposition matrix to obtain ICTs of rank $n$, weight $j$, and index $q$ $\hat{H}^{(\mathcal{V};j;q)}$ A general ICT decomposition matrix of space $\mathcal{V}$, weight $j$ and index $q$
$C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}$ A CG coefficient with input weights $l_{1}$, $l_{2}$, output weight $l_{o}$, and the corresponding indices $m_{1}$, $m_{2}$, and $m_{o}$ (the angular-momentum quantum numbers in physics)
$C^{l_{1}l_{2}l_{o}}$ A CG tensor with input weights $l_{1}$, $l_{2}$, and output weight $l_{o}$ $\hat{C}^{l_{1}l_{2}l_{o}}$ A CG matrix by merging the first two dimensions of the original CG tensor
$P^{(path)}$ A path matrix generated by $path$, e.g., $(1\to 3\to 2)$ $\hat{P}^{(path)}$ A normalized path matrix generated by $path$
$D^{(\mathcal{V}\to\mathcal{V}^{\prime})}$ A change-of-basis matrix from space $\mathcal{V}$ to $\mathcal{V}^{\prime}$ $v^{(l;p)}$ A vector on which weight $l$ and parity $p$ irreducible representation acts
${\rm Hom}_{G}$ The set of $G$-equivariant maps from one space to another ${\rm End}_{G}$ The set of $G$-equivariant maps from one space to itself
irr Returns all weights of irreducible representations in a decomposition ${\rm dim}$ The number of dimensions of a space
$(i_{1}\overset{k_{1}}{\to}i_{2}$ $\overset{k_{2}}{\to}\dots)$ A path where $k_{*}$ are bridge numbers (omitted if all being $1$) $(l,p)$ An irreducible representation space of weight $l$ and parity $p$
$\otimes$ Tensor product $\oplus$ Direct sum
$\odot$ Contraction ##\## Number of elements in a set
The following contents are organized as follows. In Section 1, we recall some basic ideas of representation theory, spherical harmonics, tensors, tensor product spaces, and we also introduce the Clebsch-Gordan coefficients and the parentage scheme. In Section LABEL:sec:problem, we formulate our problems: obtaining high-rank ICT decomposition matrices and bases for arbitrary tensor product spaces. In Section LABEL:sec:highorderict, we introduce our theory and method to obtain high-rank ICT decomposition matrices. In Section LABEL:sec:basis, we present the theory and method to obtain bases of equivariant spaces where the input and output spaces are the same. In Section LABEL:sec:extension, we extend the theory to the equivariant design spaces where the input and output spaces are different, and show the method to obtain bases for such general spaces. In Section LABEL:sec:relatedworks, we review related works on designing EGNNs and calculating ICT decomposition matrices. In Section LABEL:sec:discussion, we discuss the relation between pearce2023brauer and this work, as well as the theory for $SO(n)$, $O(n)$, and $SU(n)$. We provide concluding remarks in Section LABEL:sec:conclusion, and introduce the reproducibility details in Section LABEL:sec:repro. The acknowledgments are presented after the main text. The complexity comparison with bonvicini2024irreducible is in Appendix LABEL:sec:complexityanalysis. The visualizations of ICT decomposition matrices and equivariant bases are shown in Appendix LABEL:sec:visualization. A list of mathematical symbols is provided in Table LABEL:tab:symbols, and a mind map of this work is shown in Figure LABEL:fig:mindmap.
2 Preliminaries
In this section, we recap neccessary mathematical preliminaries.
2.1 Irreducible Representations and Spherical Spaces
We first recall some basic facts. A thorough introduction to group theory can be found in inui2012group and brocker2013representations. We first focus on the $O(3)$ group (3-D rotation and reflection), but we also extend our conclusions to $O(n)$, $SO(n)$, and $SU(n)$ in Section LABEL:sec:discussion. Consider a vector $v\in\mathcal{V}$. The rotation and reflection operations should be able to act on it. Representation theory shows that $O(3)$ can be represented by some real matrices in the Cartesian space via a group homomorphism $\rho:G\to\mathbbm{R}^{3\times 3}$, which is called the natural representation. It holds that
$\rho(g_{1}*g_{2})=\rho(g_{1})\cdot\rho(g_{2})$
for any $g_{1}$, $g_{2}\in G$, and * is the group multiplication. The group element $g$ acting on the vector $v$ can therefore be represented as a matrix-vector multiplication $\rho(g)\cdot v$. Sometimes, one may also refer to the space of $v$ as a representation space. If the results of the group action remain in the same space, i.e., $\rho(g)v\in\mathcal{V}$, this space is also called a $G$-invariant space or subspace. To represent $O(3)$, the matrix $\rho(g)$ is constrained to be orthogonal. It is important to note that the co-domain of $\rho$ can differ, meaning that it can also be represented by matrices of other dimensions. In this paper, we denote the co-domain of $\rho$ using a subscript when it should be made explicit, e.g., $\rho_{\mathbbm{R}^{3}}$. Now, it suffices to define equivariance. We claim that a linear function $f:\mathcal{V}\to\mathcal{V}^{\prime}$ is equivariant if and only if
$\rho_{\mathcal{V}^{\prime}}f(x)=f(\rho_{\mathcal{V}}(x)).$
We first focus on the simpler case where $\mathcal{V}=\mathcal{V}^{\prime}$. We will later generalize this to the case where $\mathcal{V}\neq\mathcal{V}^{\prime}$. When it is necessary to emphasize the change in space, we refer to a linear mapping $f$ as $G$-$(\mathcal{V},\mathcal{V}^{\prime})$-equivariant. We then introduce the definition of reducible and irreducible representation.
Definition 1
A group representation $\rho_{\mathcal{V}}:G\to GL(\mathcal{V})$ is said to be reducible, if there is a non-zero subspace $\mathcal{V}^{\prime}\subset\mathcal{V}$, for which we have a representation $\rho_{\mathcal{V}^{\prime}}$, such that for every $g\in G$ and $v\in\mathcal{V}^{\prime}$, it holds that
$\rho_{\mathcal{V}^{\prime}}(g)v\in\mathcal{V}^{\prime},$
where we also call $\mathcal{V}^{\prime}$ a $G$-invariant subspace. Conversely, if there is no such subspace $\mathcal{V}^{\prime}$ and the corresponding representation $\rho_{\mathcal{V}^{\prime}}$ on it for a given space $\mathcal{V}$, then we claim that $\rho_{\mathcal{V}}$ is irreducible.
If a representation is irreducible, then the vectors space it acts on is also said to be irreducible. The $3\times 3$ rotation and reflection matrices are irreducible. More generally,